OpenStudy (anonymous):

find the tangents of the following curves y^2 = 12x from (2, 7). the answers are 3x - y +1 = 0 and x- 2y + 12 = 0. helppp

OpenStudy (phi):

calculus? first, use implicit differentiation on y^2 = 12 can you do that?

OpenStudy (anonymous):

2y dy/dx = 12

OpenStudy (anonymous):

dy/dx = 12/ 2y

OpenStudy (phi):

ok, or dy/dx= 6/y

OpenStudy (anonymous):

so dy/dx = 6/7

OpenStudy (phi):

let a point on the curve be (x,y) the slope from that point to (2,7) is \[ m = \frac{y-7}{x-2} \] we want this slope to match dy/dx ( the derivative is the slope) \[ \frac{6}{y} = \frac{y-7}{x-2} \] we could solve that if we got rid of the x. the curve is y^2 = 12 x and so \[ x= \frac{y^2}{12} \] use that value for x in our equation \[ \frac{6}{y} = \frac{y-7}{\frac{y^2}{12}-2} \]

OpenStudy (phi):

can you solve for y ?

OpenStudy (anonymous):

sure, just wait sir :)

OpenStudy (anonymous):

its decimal.. not whole number or i was just too stupid solving algebra..

OpenStudy (phi):

what do you get after cross multiplying ?

OpenStudy (anonymous):

6 (y^2 / 12 -2) = y (y-7) 6y^2 / 12 -12 = y^2 - 7y 6y^2 - y^2 = -7y + 12 1/2 y^2 - 7y -12 = 0..

OpenStudy (anonymous):

am i right?? im sorry

OpenStudy (phi):

6y^2 / 12 -12 = y^2 - 7y at this step you have \[ \frac{6}{12} y^2 - 12 = y^2 -7y \\ \frac{1}{2} y^2 - 12 = y^2 -7y \] I would multiply both sides by 2 to get rid of the fraction \[ y^2 -24= 2y^2 -14y \] can you finish ?

OpenStudy (anonymous):

its still the same.. 15. 54 and -1.5

OpenStudy (anonymous):

noooo

OpenStudy (anonymous):

12 and 2

OpenStudy (phi):

\[ y^2 -24= 2y^2 -14y \\ y^2 -14y + 24 = 0 \] that factors

OpenStudy (anonymous):

its 12 and 2

OpenStudy (phi):

now that we are over that snafu... find the slope of the tangent line using m = 6/y (from the derivative way up top) we know (2,7) is a point on the line, so we can use the point slope form: y - y0= m ( x - x0) where m is the slope, and (x0,y0) is a point on the line rearrange into standard form to match your answers up top

OpenStudy (anonymous):

slope is 6/7

OpenStudy (anonymous):

y-7 = 6/7 (x-2)

OpenStudy (phi):

no, use the y values we found from \[ y^2 -14y + 24 = 0 \]

OpenStudy (anonymous):

so 6/12 or 1/2

OpenStudy (anonymous):

and 6/2 or 3

OpenStudy (phi):

yes. you will now get two different equations for the 2 lines

OpenStudy (anonymous):

got cha!! THANKS :)))))))))))